# Non-Prehensile Tool-Object Manipulation by Integrating LLM-Based Planning and Manoeuvrability-Driven Controls

### v1

@@ I Introduction
Figure 2: (a) The task environment includes a camera for real-time top-view capturing, a dual-arm robot, tool(s), and a blue block to be manipulated to the target location. (b) The architecture of our system: Unstructured data input is converted to a subtask list in the symbolic task planner with an LLM, a manoeuvrability-driven planner to compute the tool’s manoeuvrability and generate an affordance-oriented motion and path. (c) Execution process of the result given by the system: dual-arm robots take turns pushing the blue block from one side to another via collaboration.
@@
Effective tool utilisation by a robot involves primarily two aspects: (1) task planning and (2) tool movement [9, 10, 11]. Task planning is typically regarded as a cognitive high-level process in robotics, mainly used for environmental reasoning, task decomposition, allocation of action sequences, etc. [12]. However, recent trends have been pushing towards the use of LLMs to leverage the domain knowledge for semantically decomposing and planning the execution of manipulation tasks [13, 14, 15, 16]. Some examples of this directions include [14, 15], which developed an environmental feedback-based system for context-aware improvement planning. Leveraging the generative capabilities of LLMs, motion sequences can be generated for robots as demonstrated in [17, 18, 16]. The combination of traditional motion planners with LLMs has been explored in [13, 19, 20].
@@
In addition to task planning, various manipulation methodologies have been developed to model the relation between tools and objects [21]. The success of a given tool-object manipulation task largely depends on the appropriate selection of the tool. For example, robots can identify the tool type, potential uses, and contact approaches based on the tool’s geometry, see e.g., [9, 2]. In [22], tool features are learned through observation of the task’s effects and experimental validation of feature hypotheses. Affordance models are a common technique used for tool feature selection [23, 24, 25] and tool classification [26, 25, 27]. The relation between tool actions and its effects on objects is explored in [27, 28], where robots acquire affordance knowledge through predefined actions (e.g., pull, push, rotate). Recently, researchers have also explored the use of LLM in accelerating affordance learning in tool manipulation [2]. Some works have studied tool-based manipulation under constraints and from demonstrations [29, 30]. Non-prehensile object manipulation strategies have been used in [31, 32].
Although there are many studies on robotic tool use, collaborative tool-based object manipulation by dual-arm systems based on non-prehensile actions remains an underexplored problem. Notably, the challenge of applying incremental control on the stepping motion of the tool within a confined area has not been well-addressed by previous studies [21, 9, 2, 22, 29, 30, 23, 24, 26, 25, 28, 27]. This specific problem continues to present an open challenge in the field. To address this research gap, in this work, we propose a novel LLM-based manoeuvrability-driven method with the following original contributions: (1) We develop an effective model to represent the geometric-mechanical relations and manoeuvrability of tools and objects; (2) We propose a non-prehensile strategy to manoeuvre objects under different constraints with tools; (3) We evaluate the performance of the proposed methodology with real-world experiments on a dual-arm robotic system. The rest of the manuscript is organised as follows: Sec. II presents the methodology, Sec. III presents the results, Sec. IV gives final conclusions.
@@ II-A Problem Formulation
To derive our methodology, the following setup assumptions are made: (1) The manipulation motion is planar, and (2) the size of the manipulandum is not larger than any one of the segments of the tool. Throughout this paper, we denote “tool-based object manipulation” as TOM, and “tool-based object manipulation under environmental constraints” as TOME. Also, we use $\mathbf{p}^{\circ}$ to represent the 2D pose of an object $\circ$. The complete architecture of our method is depicted in Fig. 2.
@@ II-B LLM-Based High-Level Symbolic Task Planner
To obtain a valid task decomposition for a long-horizon task, the system needs to understand the requirements and generate an executable subtask list. We develop a symbolic task planner that takes natural language instructions with scene descriptions as input, and outputs a list of high-level subtasks. The list involves the tool selection/sharing between two arms, the sequence to manipulate the tool with the manipulandum, and the interaction between the two arms. The model is fine-tuned with around 20,000 lists of example data. These two data sets are generated randomly with different environment settings, such as different locations of the manipulandum, target destination, robots, and tool shapes.
@@ II-D Manoeuvrability Analysis
A tool can push the manipulandum from the side, from the tip, or other areas. However, the relative location of the manipulandum respective to the tool affects its manoeuvrability. In other words, the affordance provided by the tool is proportional to manoeuvrability. Consider using a rotating stick to push an object with its end tip. In this situation, the tool may lose contact with the manipulandum as it rolls outwards, hence, the manoeuvrability of this point is low. On the other hand, the midpoint of the stick has a high manoeuvrability, which proportionally decreases as the contact point is further away from the midpoint. We can model this behaviour with a Gaussian function, where its centre is the segment’s centre and the peak height is half the segment’s length, see Fig. 3(b). We refer to this region as an affordance area.
@@
where $n$ is the number of segments. The affordance of tool segment is quantified with the (normalised) manoeuvrability matrix: $\mathbf{M}=\hat{\mathbf{I}}/\hat{I}_{\text{max}}$, for $\hat{I}_{\text{max}}$ as the maximum value in $\hat{\mathbf{I}}$.
@@ II-E1 Initial and Final Poses
The tool’s pose corresponds to its grasping configuration, which coincides with the robot end-effector’s pose when the robot grasps the tool (see Fig. 4). We use $\mathbf{p}^{\text{tool}}$ to denote the tool’s grasping point ($x,y$ coordinates) when it has not come in contact with the object. To construct a trajectory for tool-based object transport, we need to find out the tool’s desired initial and final poses for the subtask “$\mathtt{interact}$”. We first define these poses (which include the orientation) of the chosen tool as $\mathbf{p}^{\text{int}}_{\text{start}}$ and $\mathbf{p}^{\text{int}}_{\text{end}}$ respectively.
@@
The motion trajectory of a tool, moving along the z-axis of the object’s centre without displacing it can be described as a circular trajectory with the centre $\mathbf{p}^{\text{obj}}$ and radius $r$, where $r=\|\mathbf{p}_{\ast}-\mathbf{p}^{\text{tool}}\|$. We represent the trajectories for the initial and final configurations as $\mathbf{C}_{\text{start}}$ and $\mathbf{C}_{\text{end}}$ (see Fig. 4(a)).
@@ II-F Application with Environmental Constraints
Consider the manipulandum is tightly confined within a concave-shaped wall, as shown in Fig. 5(a), with an unknown exit and assume that the tool can enter the constrained area. To move the manipulandum out from the bounded area with small movement space, we determine the direction from the manipulandum to the exit by considering the overall affordance of the wall boundary. We denote this direction vector as $\mathbf{v}^{\text{exit}}$, and its magnitude is defined as the minimum travel distance for the manipulandum. We consider the inner edge of the wall as a segment $\mathbf{s}^{\text{wall}}_{i}$ where $i=\{1,\dots,n^{\text{wall}}\}$ and $n^{\text{wall}}$ is the number of the wall segment. The affordance of a wall is passively provided and is defined as $\mathbf{a}^{\text{wall}}_{i}$ with the model shown in Sec. II-C. The passive affordance vector is the normal vector of $\mathbf{s}^{\text{wall}}_{i}$ located in the middle with the direction pointing towards the constrained area. Its magnitude is scaled to half of $\mathbf{s}^{\text{wall}}_{i}$ as the block is generally not receiving any affordance from a wall segment based on our visual affordance model. The moving direction for the manipulandum to the exit can be obtained by the following equation:
@@
To move in the limited area while interacting with the block, we employ a stepping approach to manipulate the block in the confined area. As the possible movement area is small and highly restricted, an incremental pulsing motion is adopted to make small adjustments with high accuracy motion control to the tool and the manipulandum. Inspired by the animal manipulation study in [4] (where a crow uses a tool to get the food from the box slot by rotating and dragging the tool outwards), we adopt a similar approach to retrieve the object from confined spaces. This strategy continuously alternates between “repositioning” the tool and incremental “rotation-dragging” the object towards the exit until it can be fully extracted. This strategy is illustrated in Fig. 5.
@@
We define $\tau$ as an action step variable and is incremented by $1$ if an action (reposition/rotation-dragging) is fulfilled (i.e. $\tau=0,1,2,\dots$). To control the change of action, a step function (denoted as $u(\tau)$) is implemented as a trigger with the step variable $\tau$. This kind of non-prehensile crow-inspired behaviour can be unified and modelled as:
@@
$f(\phi_{\tau+1})=\left\{\begin{aligned} &-\angle{\text{obj}}-\angle{\text{rot}}\quad\text{if direction is anti-clockwise}\\ &-\phi_{\tau}+\pi-\angle{\text{obj}}-\angle{\text{rot}}\quad\text{otherwise}\end{aligned}\right.$ (9)
@@ III Results
To validate our methodology in terms of accuracy and robustness, we have conducted around 200 experiments in a dual-arm robot system. In the experiment, two sets of UR-3 robotic arms are used and GPT 3.5 is implemented for task decomposition. Three types of tools are selected which are a stick, an L-shaped hook, and a Y-shaped hook (see Fig.1). Different tool combinations are evaluated with diverse movement directions and tasks. Various masses of the block is tested and validated. Since this is a vision-based controller, the mass of the block does not significantly affect the results. Therefore, the block’s mass is omitted from this section. A RealSense D415 captures the images of the whole process. Data is passed to a Linux-based computer with the Robot Operating System (ROS) for image process and robot control. Aruco code is used for providing accurate pose tracing in real time.
@@ III-A Single-Arm Robot with a Single Tool
We first evaluate the task decomposition performance of LLM. For that, a tool and a blue block are placed on the table with the target given as shown in Fig. 6. The task is to manipulate the block within a close distance, which is sufficient for a single-arm robot. The embedded information which contains the task, the environment and the geometry of the tool is passed to the LLM. In the experiment shown in Fig. 6(a), the robot executes the subtasks generated by the high-level symbolic task planner which include: $\mathtt{grasp}(right,hook);$ $\mathtt{approach}(right,hook,block);$ $\mathtt{interact}(right,hook,block,target);\mathtt{release}(right,hook)$, where the right arm first moves and grasps the hook, then moves the block to the target, and lastly releases the tool back to its original place. In a non-single tool scenario, where two tools are available on the desk as shown in Fig. 6(c), the task planner selects the nearest tool based on the embedded information to push the block towards the target. The experiment showcases the application of the proposed affordance and manoeuvrability model in locating the highest manoeuvrability region for block transportation. During the manipulation stage, the block is kept within the highest manoeuvrability region (indicated with a red circle in Fig. 6) to receive affordance effectively from the tool. The minimisation of the error between the $\mathbf{p}^{\text{obj}}$ and the $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 7. These results corroborate that the proposed method can be used to actively drive a robot to manipulate an object via a tool.
@@ III-B Dual-Arm Robot with Long-Horizon Task
We then evaluate the long-horizon task performance where the block has to travel from far right to far left, far right/left to top right/left, and vice versa. The long-horizon task is evaluated with multiple tool combinations. The system observes and generates a collaborative motion plan. In the experiment shown in Fig. 8(a), the right and left arms pick up the stick and the hook respectively. The right arm uses the stick to push the block to the left side, allowing the left arm to continue the task. The robot leverages the advantage of the hook to drag the block closer to its working area and push the block to the desired location. In Fig. 8(b), the right and left arms grasped the Y-shaped tool and the stick respectively. The right arm uses the tool to pass the block to the left. The left arm uses the stick to push the block to the target location.
Figure 6: Single-arm robot with a single tool: moving the block from (a) right to left with a hook, (b) right to left with a stick, (c) bottom to top with a Y-shaped tool. The trajectory of the block is reflected in the red line. The highest manoeuvrability point is indicated with a red circle.
@@
Figure 8: Long-horizon task: moving the block from (a) far right to far left with a hook and a stick, (b) far top right to far left with a stick and a Y-shaped tool, (c) far left to far right with a hook; and (d)–(f) exit from a confined area with a stepping controller. The block trajectory is reflected in pink and the target is labelled with a blue square.
@@
The long-horizon task performance is evaluated with the tool-sharing ability. Assuming there is only one tool available, it has to be shared among the dual-arm robot. Fig. 8(c) demonstrates the tool is passed to another arm once the block is pushed to the middle of the table. The block is moved accurately to the target with motion-decomposed: ‘$\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{pass};$ $\mathtt{release};$ $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{release}$’ where the left arm releases the tool once it is done and the right picks up the tool to continue moving the block. Though the hook is in a two-link geometry, the pushing is afforded by the right side of the tool (a single segment) with the highest manoeuvrability region.
The minimisation of the error between $\mathbf{p}^{\text{obj}}$ and $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 9. Similar to the single-arm robot with a single tool experiment, this long-horizon task experiment also demonstrates the robustness of the proposed methodology such that the tasks are successfully decomposed into multiple collaborative subtasks, and the highest manoeuvrability region of the tool is leveraged in block manipulation.

### v2

@@ I Introduction
Figure 2: (a) The task environment includes a camera for real-time top-view capturing, a dual-arm robot, tool(s), and a blue manipulandum to be manipulated to the target location. (b) The architecture of our system: Unstructured data input is converted to a subtask list in the symbolic task planner with an LLM, a manoeuvrability-driven planner to compute the tool’s manoeuvrability and generate an affordance-oriented motion and path. (c) Execution process of the result given by the system: dual-arm robots take turns pushing the blue manipulandum from one side to another via collaboration.
I-A Related Works
@@
Effective tool utilisation by a robot involves primarily two aspects: (1) task planning and (2) tool movement [9, 10, 11]. Task planning is typically regarded as a cognitive high-level process in robotics, mainly used for environmental reasoning, task decomposition, allocation of action sequences, etc. [12]. Task can be decomposed with the integration of learning-based approaches, particularly through the use of reinforcement learning techniques to optimize task planning [13, 14, 15]. Studies have also highlighted the effectiveness of rule-based planning methods, which incorporate predefined heuristics and logical rules to enhance the efficiency of task decomposition in structured environments [16, 17, 18, 19]. While rule-based planning is effective for well-defined problems, it can struggle with complex, dynamic environments where the number of rules may become unmanageable. However, recent trends have been pushing towards the use of LLMs to leverage the domain knowledge for semantically decomposing and planning the execution of manipulation tasks [20, 21, 22, 23, 24, 25]. Some examples of this direction include [21, 22], which developed an environmental feedback-based system for context-aware improvement planning. Leveraging the generative capabilities of LLMs, motion sequences can be generated for robots as demonstrated in [26, 27, 23]. The combination of traditional motion planners with LLMs has been explored in [20, 28, 29]. Domain knowledge can be integrated with LLMs to generate a list of motions for navigating a robot in an apartment, as demonstrated in [25]. However, the focus primarily remains on independent motions. Motivated by [25], we further consider the dependent motion among arms and tools.
@@
Transitioning from the critical role of task planning, it is evident that effective tool use is inherently tied to understanding the relationship between tools and objects. Indeed, the success of a given tool-object manipulation task largely depends on the appropriate selection of the tool, which necessitates a nuanced comprehension of how tools interact with various objects in their environment [30]. For example, robots can identify the tool type, potential uses, and contact approaches based on the tool’s geometry, see e.g., [9, 2]. In [31], tool features are learned through observation of the task’s effects and experimental validation of feature hypotheses. Affordance models are a common technique used for tool feature selection [32, 33, 34] and tool classification [35, 34, 36]. The relation between tool actions and its effects on objects is explored in [36, 37], where robots acquire affordance knowledge through predefined actions (e.g., pull, push, rotate). Recently, researchers have also explored the use of LLM in accelerating affordance learning in tool manipulation [2]. Some works have studied tool-based manipulation under constraints and from demonstrations [38, 39]. Non-prehensile object manipulation strategies have been used in [40, 41].
Building on this foundation of understanding tool-object interactions, it is important to highlight that, despite the advancements in robotic tool use, collaborative tool-based object manipulation by dual-arm systems based on non-prehensile actions remains an underexplored problem. Notably, the challenge of applying incremental control on the stepping motion of the tool within a confined area has not been well-addressed by previous studies [30, 9, 2, 31, 38, 39, 32, 33, 35, 34, 37, 36]. Furthermore, most studies have primarily focused on task decomposition for simple object manipulation using LLMs, with tool manipulation being rarely addressed. Dual-arm collaborative manipulation utilizing non-prehensile tools represents a promising area for further exploration. In other words, the integration of LLMs in tool-object manipulation with dual-arm robots remains underexplored. This specific challenge continues to present an open opportunity in the field.
I-B Contributions
To address this research gap, in this work, we propose a novel LLM-based manoeuvrability-driven method with the following original contributions: (1) We develop an effective model to represent the geometric-mechanical relations and manoeuvrability of tools and objects; (2) We propose a non-prehensile strategy to manoeuvre objects under different constraints with tools; (3) We evaluate the performance of the proposed methodology with real-world experiments on a dual-arm robotic system. Our work uniquely integrates LLMs to enhance tool-object interactions, enabling robots to interpret and execute complex non-prehensile manipulation tasks through natural language instructions. This integration allows for dynamic adaptation to various situations and fosters intuitive human-robot collaboration, significantly improving the effectiveness of dual-arm tool-object manipulation.
The rest of the manuscript is organised as follows: Sec. II presents the methodology, Sec. III presents the results, Sec. IV gives final conclusions.
@@ II-A Problem Formulation
To derive our methodology, the following setup assumptions are made: (1) The manipulation motion is planar, (2) the size of the manipulandum is not larger than any one of the segments of the tool, and (3) the manipulandum has a simple, regular geometric shape, such as circular or hexagonal. Throughout this paper, “tool-based object manipulation” is denoted as TOM, and “tool-based object manipulation under environmental constraints” is denoted as TOME. Also, $\mathbf{p}^{\circ}$ represents the 2D pose of an object $\circ$. The complete architecture of our method is depicted in Fig. 2.
@@ II-B LLM-Based High-Level Symbolic Task Planner
To obtain a valid task decomposition for a long-horizon task, the system needs to understand the requirements and generate an executable subtask list. We develop a symbolic task planner that takes natural language instructions with scene descriptions as input, and outputs a list of high-level subtasks. The list involves the tool selection/sharing between two arms, the sequence to manipulate the tool with the manipulandum, and the interaction between the two arms. The model is fine-tuned using approximately 20,000 example data lists, specifically tailored for our non-prehensile tool object manipulation scenario. During the fine-tuning stage, we utilized a program to create 20,000 distinct environmental setups by randomly varying the poses of the robot, tool, block, and target within a finite combination space. This approach allows task decomposition to be framed as a classification problem, enabling the LLM to effectively correlate each setup with a specific list of motion functions, thereby enhancing its ability to predict expected outcomes based on prompt patterns rather than producing hallucinations.
@@ II-D Manoeuvrability Analysis
A tool can push the manipulandum from the side, from the tip, or other areas. However, the relative location of the manipulandum respective to the tool affects its manoeuvrability. In other words, the affordance provided by the tool is proportional to manoeuvrability. Consider using a rotating stick to push an object with its end tip. In this situation, the tool may lose contact with the manipulandum as it rolls outwards, hence, the manoeuvrability of this point is low. On the other hand, the midpoint of the stick has a high manoeuvrability, which proportionally decreases as the contact point is further away from the midpoint. This behaviour can be modelled with a Gaussian function, where its centre is the segment’s centre and the peak height is half the segment’s length, see Fig. 3(b). We refer to this region as an affordance area.
@@
where $n$ is the number of segments. The affordance of the tool segment is quantified with the (normalised) manoeuvrability matrix: $\mathbf{M}=\hat{\mathbf{I}}/\hat{I}_{\text{max}}$, for $\hat{I}_{\text{max}}$ as the maximum value in $\hat{\mathbf{I}}$.
@@ II-E1 Initial and Final Poses
The tool’s pose corresponds to its grasping configuration, which coincides with the robot end-effector’s pose when the robot grasps the tool (see Fig. 4). $\mathbf{p}^{\text{tool}}$ denotes the tool’s grasping point ($x,y$ coordinates) when it has not come in contact with the object. To construct a trajectory for tool-based object transport, we need to find out the tool’s desired initial and final poses for the subtask “$\mathtt{interact}$”. We first define these poses (which include the orientation) of the chosen tool as $\mathbf{p}^{\text{int}}_{\text{start}}$ and $\mathbf{p}^{\text{int}}_{\text{end}}$ respectively.
@@
The motion trajectory of a tool, moving along the z-axis of the object’s centre without displacing it can be described as a circular trajectory with the centre $\mathbf{p}^{\text{obj}}$ and radius $r$, where $r=\|\mathbf{p}_{\ast}-\mathbf{p}^{\text{tool}}\|$. The trajectories for the initial and final configurations are represented as $\mathbf{C}_{\text{start}}$ and $\mathbf{C}_{\text{end}}$ (see Fig. 4(a)).
@@ II-F Application with Environmental Constraints
Consider the manipulandum is tightly confined within a concave-shaped wall, as shown in Fig. 5(a), with an unknown exit and assume that the tool can enter the constrained area. To move the manipulandum out from the bounded area with small movement space, we determine the direction from the manipulandum to the exit by considering the overall affordance of the wall boundary. We denote this direction vector as $\mathbf{v}^{\text{exit}}$, and its magnitude is defined as the minimum travel distance for the manipulandum. Consider the inner edge of the wall as a segment $\mathbf{s}^{\text{wall}}_{i}$ where $i=\{1,\dots,n^{\text{wall}}\}$ and $n^{\text{wall}}$ is the number of the wall segment. The affordance of a wall is passively provided and is defined as $\mathbf{a}^{\text{wall}}_{i}$ with the model shown in Sec. II-C. The passive affordance vector is the normal vector of $\mathbf{s}^{\text{wall}}_{i}$ located in the middle with the direction pointing towards the constrained area. Its magnitude is scaled to half of $\mathbf{s}^{\text{wall}}_{i}$ as the manipulandum is generally not receiving any affordance from a wall segment based on our visual affordance model. The moving direction for the manipulandum to the exit can be obtained by the following equation:
@@
To move in the limited area while interacting with the manipulandum, we employ a stepping approach to manipulate the manipulandum in the confined area. As the possible movement area is small and highly restricted, an incremental pulsing motion is adopted to make small adjustments with high accuracy motion control to the tool and the manipulandum. Inspired by the animal manipulation study in [4] (where a crow uses a tool to get the food from the box slot by rotating and dragging the tool outwards), we adopt a similar approach to retrieve the object from confined spaces. This strategy continuously alternates between “repositioning” the tool and incremental “rotation-dragging” the object towards the exit until it can be fully extracted as depicted in Fig. 5.
@@
$\tau$ is an action step variable and is incremented by $1$ if an action (reposition/rotation-dragging) is fulfilled (i.e. $\tau=0,1,2,\dots$). To control the change of action, a step function (denoted as $u(\tau)$) is implemented as a trigger with the step variable $\tau$. This kind of non-prehensile crow-inspired behaviour can be unified and modelled as:
@@
$f(\phi_{\tau+1})=\left\{\begin{aligned} &-\angle{\text{obj}}-\angle{\text{rot}},~{}~{}\text{if direction is anti-clockwise}\\ &-\phi_{\tau}+\pi-\angle{\text{obj}}-\angle{\text{rot}},~{}~{}\text{otherwise}\end{aligned}\right.$ (9)
@@ III Results
To validate our methodology in terms of accuracy and robustness, we have conducted around 200 experiments in a dual-arm robot system. In the experiment, two sets of UR-3 robotic arms are used and GPT 3.5 is implemented for task decomposition. Three types of tools are selected which are a stick, an L-shaped hook, and a Y-shaped hook (see Fig.1). Different tool combinations are evaluated with diverse movement directions and tasks. Various masses of the manipulandum are tested and validated. Since this is a vision-based controller, the mass of the manipulandum does not significantly affect the results. Therefore, the manipulandum’s mass is omitted from this section. A RealSense D415 captures the images of the whole process. Data is passed to a Linux-based computer with the Robot Operating System (ROS) for image process and robot control. Aruco code is used for providing accurate pose tracing in real time. The average inference time is approximately 0.158 seconds for tool analysis and around 1.51 seconds for LLM processing. Since these operations are completed before the robot begins its movements, their effect on overall system responsiveness remains minimal.
@@ III-A Single-Arm Robot with a Single Tool
We first evaluate the task decomposition performance of LLM. For that, a tool and a blue manipulandum are placed on the table with the target given as shown in Fig. 6. The task is to manipulate the manipulandum within a close distance, which is sufficient for a single-arm robot. The embedded information which contains the task, the environment and the geometry of the tool is passed to the LLM. In the experiment shown in Fig. 6(a), the robot executes the subtasks generated by the high-level symbolic task planner which include: $\mathtt{grasp}(right,hook);$ $\mathtt{approach}(right,hook,block);$ $\mathtt{interact}(right,hook,block,target);\mathtt{release}(right,hook)$, where the right arm first moves and grasps the hook, then moves the block to the target, and lastly releases the tool back to its original place. In a non-single tool scenario, where two tools are available on the desk as shown in Fig. 6(c), the task planner selects the nearest tool based on the embedded information to push the block towards the target. The experiment showcases the application of the proposed affordance and manoeuvrability model in locating the highest manoeuvrability region for manipulandum transportation. During the manipulation stage, the manipulandum is kept within the highest manoeuvrability region (indicated with a red circle in Fig. 6) to receive affordance effectively from the tool. The minimisation of the error between the $\mathbf{p}^{\text{obj}}$ and the $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 7. These results corroborate that the proposed method can be used to actively drive a robot to manipulate an object via a tool.
@@ III-B Dual-Arm Robot with Long-Horizon Task
We then evaluate the long-horizon task performance where the manipulandum has to travel from far right to far left, far right/left to top right/left, and vice versa. The long-horizon task is evaluated with multiple tool combinations. The system observes and generates a collaborative motion plan. In the experiment shown in Fig. 8(a), the right and left arms pick up the stick and the hook respectively. The right arm uses the stick to push the manipulandum to the left side, allowing the left arm to continue the task. The robot leverages the advantage of the hook to drag the manipulandum closer to its working area and push the manipulandum to the desired location. In Fig. 8(b), the right and left arms grasped the Y-shaped tool and the stick respectively. The right arm uses the tool to pass the manipulandum to the left. The left arm uses the stick to push the manipulandum to the target location.
Figure 6: Single-arm robot with a single tool: moving the manipulandum (a) right to left with a hook, (b) right to left with a stick, and (c) bottom to top with a Y-shaped tool. The red line shows the manipulandum’s trajectory, while the red circle indicates the highest maneuverability point.
@@
Figure 8: Long-horizon task: moving the manipulandum from (a) far right to far left with a hook and a stick, (b) far top right to far left with a stick and a Y-shaped tool, (c) far left to far right with a hook; and (d)–(f) exit from a confined area with a stepping controller. The manipulandum trajectory is reflected in pink and the target is labelled with a blue square.
@@
The long-horizon task performance is evaluated with the tool-sharing ability. Assuming there is only one tool available, it has to be shared among the dual-arm robot. Fig. 8(c) demonstrates the tool is passed to another arm once the manipulandum is pushed to the middle of the table. The manipulandum is moved accurately to the target with motion-decomposed: ‘$\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{pass};$ $\mathtt{release};$ $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{release}$’ where the left arm releases the tool once it is done and the right picks up the tool to continue moving the manipulandum. Though the hook is in a two-link geometry, the pushing is afforded by the right side of the tool (a single segment) with the highest manoeuvrability region.
The minimisation of the error between $\mathbf{p}^{\text{obj}}$ and $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 9. Similar to the single-arm robot with a single tool experiment, this long-horizon task also demonstrates the robustness of the proposed methodology such that the tasks are successfully decomposed into multiple collaborative subtasks, and the highest manoeuvrability region of the tool is leveraged in manipulandum manipulation.